# Federated Large Language Models: Current Progress and Future Directions

### v1

@@ Abstract
Large language models are rapidly gaining popularity and have been widely adopted in real-world applications. While the quality of training data is essential, privacy concerns arise during data collection. Federated learning offers a solution by allowing multiple clients to collaboratively train LLMs without sharing local data. However, FL introduces new challenges, such as model convergence issues due to heterogeneous data and high communication costs. A comprehensive study is required to address these challenges and guide future research. This paper surveys Federated learning for LLMs (FedLLM), highlighting recent advances and future directions. We focus on two key aspects: fine-tuning and prompt learning in a federated setting, discussing existing work and associated research challenges. We finally propose potential directions for federated LLMs, including pre-training and LLMs for federated learning.
@@ 1 Introduction
[6] [88] [72] [106] Ours
@@
Efficiency
Potential Directions
Table 1: Comparison with Former Surveys.
The integration of FL with LLMs introduces unique challenges and opportunities. The sheer size of LLMs and the complexity of their training processes require careful consideration of communication overhead, computational efficiency, and convergence stability. Researchers have started to address these issues, leading to a growing body of literature that explores various aspects of Federated Learning for LLMs (FedLLM). Several surveys have attempted to capture the evolving landscape of FedLLM research. For instance, [6] and [88] focus on privacy protection mechanisms within FL frameworks for LLMs.  [72] take a different approach by concentrating on communication and computation efficiency. They analyze strategies to reduce the communication overhead between clients and the central server, such as model compression, sparsification, and quantization.  [106] offers insights into the future of foundation models, including LLMs, but lacks a comprehensive review of recent papers on federated learning.
@@ 2 Federated Fine-Tuning
Fine-tuning is one of the most important directions of FedLLMs. Traditional topics in FL are still relevant but introduce new challenges in the context of FedLLMs, as shown in Table 2. Efficiency plays a key role in reducing the high cost of training. Personalization enables the creation of AI assistants tailored to individual users. Privacy and security become even more critical in the context of LLMs, given the sensitive nature of the data involved. Instruction-tuning and value alignment help enhance model accuracy and mitigate risks in applying LLMs.
@@
FedDAT [8], RaFFM [87], FedKC [61],
 [62], FEDSA [97]
@@
FedLoRA [86], HETLORA [12], FlexLoRA [2],
FedPCL [56], FedBRB [81]
Privacy and Security Security and Robustness  [41],  [100][73],  [25]
Privacy FedPIT [96], FFA-LoRA [55]
Attacks  [73],  [24], Decepticon [18],  [33]
Defense FEDML-HE [28],  [88], FedBiOT [74]
Efficiency Training Efficiency Grouper [4], FedYolo [93], FedTune [10]
Communication Efficiency CEFHRI [30], FedKSeed [46], FedRDMA [94]
@@
 [52], FedPETuning [95],  [63],
SLoRA [1], FFA-LoRA [54],
FlexLoRA [2], LP-FL [27], FedBiOT [75]
Frameworks Cross-Silo FedRDMA [94], CROSSLM [17]
Cross-Device  [60], FwdLLM [79],  [71]
Decentralized Training OpenFedLLM [85],  [57],  [89]
Blackbox and Transfer Learning Fed-BBPL  [35], ZooPFL  [43],  Kang et al. [29]
Instruction Tuning FederatedScope-LLM [31], FedIT  [90]
Evaluation Datasets FederatedScope-LLM [31], OpenFedLLM [85]
Benchmarks FedLLM-Bench  [84], Fedmlsecurity [26], Profit [14]
Convergence Analysis FedPEAT [13], FedMeZO [36]
Table 2: Overview of Topics in Federated Fine-Tuning
@@ 2.1 Heterogeneity
Data Heterogeneity To address data heterogeneity for various vision-language tasks in federated multimodal fine-tuning, FedDAT [8] proposes a Dual-Adapter Teacher to regularize client’s local updates and Mutual Knowledge Distillation for knowledge transfer. RaFFM [87] proposes a novel model compression method to accommodate heterogeneous training data. [61] focuses on addressing the challenges of data heterogeneity in multilingual federated natural language understanding by proposing a FedKC module that exchanges knowledge across clients without sharing raw data. [62] proposes a distillation method from heterogeneous tag sets to achieve better domain adaptation. In addition to transferring multimodal knowledge in FL, M2FEDSA [97] proposes split learning to realize modularized decomposition and enables unimodal local models to be complementarily aggregated.
Personalization and Model Heterogeneity Heterogeneity FedLoRA [86] designs a homogeneous adapter to accommodate federated client’s heterogeneous local model training, which can be aggregated on the fine-tuning server. HETLORA [12] instead uses heterogeneous LoRA designs across devices to reduce training imbalanced problems, and sparsity-weighted aggregation is developed on the server. FlexLoRA [2] also enables heterogeneous LoRA ranks, which are aggregated on the server using Singular Value Decomposition. FedPCL [56] is proposed to federally fine-tune client-specific and class-relevant information with homogeneous local models as prototypes. By prototype-wise contrastive learning, the aggregated model can preserve personalized model representation, while achieving similar performance with a large-scale centrally trained model. FedBRB [81] also proposes a small-to-large FL framework, where small local models are fine-tuned as blocks of a large global model.
@@ 2.2 Privacy and Security
Security and Robustness [41] comprehensively discussed potential vulnerabilities and essential components in decentralized training for LLMs in terms of data and models. [100] studied a comprehensive literature survey about privacy-preserving in federated training of foundation models. [73] investigated the vulnerability of federated foundation models under adversarial threats. In addition, a novel attack strategy is developed to target the safety issues of foundation models in the setting of FL.
Privacy To enhance privacy protection and improve model performance, FedPIT [96] proposed to use LLM’s in-context learning abilities to augment local synthetic datasets. FFA-LoRA [55] fixes the randomly initialized non-zero matrices and only fine-tune the zero-initialized matrices in LoRA to solve learning inefficiency in privacy-preserved FL.
Attacks Several papers also discuss the potential adversarial attack strategies and situations in federated LLMs fine-tuning. Wu et al.  [73] investigated the vulnerability in federated fine-tuning under adversarial threats. Experiments are conducted in both image and text domains, where this work discovered the high susceptibility of federated fine-tuning to adversarial attacks. Attackers can also recover private text in federated learning of language models [24] with high fidelity. Further experiments show that if in fine-tuning stage, the pre-trained language model’s word embedding remains unchanged, such attacks can be less effective. Decepticon [18] also proposes a latent vector attack, which can discover private user text with malicious parameter vectors. In heterogeneous federated learning, Li et al.  [33] also develop a novel backdoor attack mechanism by mimicking normal client behavior.
Defense Defense for privacy-preserving and malicious attacks can be beneficial and challenging for LLMs, due to the efficiency-effectiveness balance for large-scale models. FEDML-HE [28] proposes a homomorphic encryption algorithm that selectively encrypts sensitive parameters and balances security and efficiency. The optimized federated fine-tuning method can significantly reduce the computation overhead. [88] more comprehensively studied the federated foundation model in privacy-preserving pre-training, fine-tuning, and application. In addition, intellectual property protection is also crucial in federated LLM fine-tuning, where FedBiOT [74] creates a bi-level optimization problem to ensure the emulator, trained on a public dataset by the LLM owner supporting adaptors in fine-tuning on clients’ private datasets, which works effectively regardless of distribution drift between the datasets.
@@ 2.3 Efficiency
Training Efficiency In LLM-based federated learning, optimizing training efficiency is crucial for minimizing latency and computational overhead. This ensures rapid and effective fine-tuning across diverse and distributed private datasets. Dataset Grouper [4], creates large-scale federated datasets to address the data efficiency problem in federated LLM fine-tuning, which support flexible and scalable federated learning simulations for training language models with billions of parameters. FedYolo [93] explores possibilities in on-device training of large language models, emphasizing the benefits of model scale and modularity to enhance heterogeneity robustness, reduce communication costs, and prevent catastrophic forgetting while enabling clients to adapt to diverse tasks with a single general-purpose model. FedTune [10] benchmarks three fine-tuning methods (e.g., modifying the input, adding extra modules, and adjusting the backbone) and conducts empirical experiments to show the comprehensive comparison between federated learning and ordinary fine-tuning.
@@
Communication Efficiency Managing communication costs while ensuring data privacy is paramount in federated fine-tuning for large language models. Recent research has explored innovative solutions to reduce communication overhead while maintaining or enhancing the performance of federated systems. The integration of federated learning in HRI and robots benefits from data privacy but suffers from communication inefficiency. Khalid et al. [30] propose a communication-efficient FL framework for HRI (CEFHRI), leveraging pre-trained models and a trainable spatiotemporal adapter for video understanding tasks. FedKSeed [46], explores federated full-parameter fine-tuning, which can be communication inefficient, in LLMs, which utilizes zeroth-order optimization with finite random seeds, significantly reducing transmission requirements to a few thousand bytes. FedRDMA [94], a cross-silo FL system, integrates RDMA into the FL communication protocol to enhance efficiency, by dividing the updated model into chunks and optimizing RDMA-based communication.
@@
Parameter Efficiency Several existing papers have explored parameter-efficient learning methods for fine-tuning large language models (LLMs) in federated learning (FL) [75]. These methods address the communication, computation, and storage constraints of edge devices, making federated training more efficient. [52] systematically investigate ways to reduce communication in federated fine-tuning, while [95] study four PETuning methods and propose a benchmark to defend against privacy attacks. [63] introduce a gradient-free prompt tuning method for better performance in few-shot settings. In heterogeneous FL environments, several LoRA-based methods have been developed: SLoRA [1] enables fine-tuning with sparse parameter masks and SVD-decomposed LoRA blocks, FFA-LoRA [54] mitigates LoRA instability due to data heterogeneity, and FlexLoRA [2] adjusts LoRA ranks dynamically. LP-FL [27] leverages few-shot prompt learning and soft labeling for efficient communication.
@@ 2.4 Frameworks
Cross-Silo Federated learning (FL) in cross-silo settings presents unique challenges, particularly in communication efficiency and collaborative learning between different model sizes. FedRDMA [94] is a communication-efficient FL system integrating Remote Direct Memory Access (RDMA) to alleviate the communication overheads exacerbated by LLMs. FedRDMA significantly improves communication efficiency over traditional TCP/IP-based systems by segmenting the updated model and employing optimization techniques. CROSSLM [17] uses smaller language models (SLMs) trained on private task-specific data to alleviate privacy concerns, which can further augment the learning of LLMs. This mutual enhancement approach allows both SLMs and LLMs to improve their performance while preserving LLMs’ generalization capabilities.
@@
Cross-Device Different from cross-silo settings, normally cross-device federated learning involves a large number of small-scale end devices, where fine-tuning LLMs can be challenging. A systematic study [60] has used large-scale public data and LLMs to explore distillation methods in training large-scale parallel devices to improve privacy-utility tradeoff. To address the challenge of resource constraints on mobile devices, FwdLLM [79] proposes a backpropagation-free training protocol that enhances memory and time efficiency in federated learning of LLMs, facilitating the deployment of billion-parameter LLMs on commercial mobile devices. Woisetschläger et al. [71] focus on energy efficiency in federated learning of LLMs on edge devices. The FLAN-T5 foundation models are evaluated end-to-end for fine-tuning with a real-time metric of energy efficiency to assess the computational efficiency.
@@
Decentralized Training The extensive computational resources required for training and deploying LLMs have become a significant challenge. Recent research has explored decentralized FL as a solution, leveraging underutilized consumer-level GPUs and private data to make LLM training more accessible and privacy-preserving. OpenFedLLM [85] proposes a complete pipeline for training LLMs on the underutilized distributed private data via FL. Specifically, to leverage consumer-level GPUs for LLM training, Tang et al. [57] proposed a decentralized system that includes a dynamic broker with a backup pool to handle the variability and heterogeneity of devices. Intermediate representation and execution planes are abstracted to ensure compatibility across different devices and deep learning frameworks. Yuan et al. [89] propose a scheduling algorithm that efficiently allocates computational tasks across decentralized GPUs connected by slow networks, which enables decentralized training among heterogeneous and lower-bandwidth interconnected devices.
@@
Blackbox and Transfer Learning When dealing with black-box foundation models, gradients are often unavailable, where zeroth-order optimization techniques are explored as a means to update models without requiring gradient access. Fed-BBPL  [35] optimizes prompt generators in federated learning settings without directly accessing the foundation models’ parameters for text and image classification tasks. ZooPFL  [43] presents a more general approach making it suitable for more complex and diverse FL tasks. Kang et al. [29] instead focus on transfer learning that enables clients to benefit from general knowledge learned in federated settings and also transfer domain-specific knowledge back to improve the federated learning.
@@
Instruction Tuning FederatedScope-LLM [31] includes multiple efficient federated fine-tuning algorithms to reduce communication and computation costs, where clients may or may not have access to the full model parameters. FedIT  [90] designs FL approaches to incorporate diverse user instructions on local devices, which might be data sensitive.

### v2

@@ Abstract
Large language models are rapidly gaining popularity and have been widely adopted in real-world applications. While the quality of training data is essential, privacy concerns arise during data collection. Federated learning offers a solution by allowing multiple clients to collaboratively train LLMs without sharing local data. However, FL introduces new challenges, such as model convergence issues due to heterogeneous data and high communication costs. A comprehensive study is required to address these challenges and guide future research. This paper surveys Federated learning for LLMs (FedLLM), highlighting recent advances and future directions. We focus on two key aspects: fine-tuning and prompt learning in a federated setting, discussing existing work and associated research challenges. We finally propose potential directions for federated LLMs, including pre-training, federated agents, and LLMs for federated learning.
@@ 1 Introduction
Table 1: Comparison with Existing Surveys on Federated Large Language Models.
[chen2023federated] [yu2023federated] [woisetschlager2024survey] [zhuang2023foundation] Ours
@@
Communication & Computation Efficiency
Prompt Tuning & PEFT Methods
Real World Applications
Emerging Directions (Agents / Pre-training)
The integration of FL with LLMs introduces unique challenges and opportunities. The sheer size of LLMs and the complexity of their training processes require careful consideration of communication overhead, computational efficiency, and convergence stability. Researchers have started to address these issues, leading to a growing body of literature that explores various aspects of Federated Learning for LLMs (FedLLM). Several surveys have attempted to capture the evolving landscape of FedLLM research. For instance, [chen2023federated] and [yu2023federated] focus on privacy protection mechanisms within FL frameworks for LLMs.  [woisetschlager2024survey] take a different approach by concentrating on communication and computation efficiency. They analyze strategies to reduce the communication overhead between clients and the central server, such as model compression, sparsification, and quantization.  [zhuang2023foundation] offers insights into the future of foundation models, including LLMs, but lacks a comprehensive review of recent papers on federated learning.
@@ 2 Federated Fine-Tuning
Fine-tuning is one of the most important directions of FedLLMs. Traditional topics in FL are still relevant but introduce new challenges in the context of FedLLMs, as shown in Table 2. Efficiency plays a key role in reducing the high cost of training. Personalization enables the creation of AI assistants tailored to individual users. Privacy and security become more critical in the context of LLMs, given the sensitive nature of the data involved. Instruction-tuning and value alignment enhance model accuracy and mitigate risks in applying LLMs.
Table 2: Overview of Topics in Federated Fine-Tuning.
@@
FedDAT [chen2023feddat], RaFFM [yu2023bridging], FedKC [wang2022fedkc],
 [wang2023federated], FEDSA [zhangenhancing]
@@
FedLoRA [yi2023fedlora], HETLORA [cho2024heterogeneous], FlexLoRA [bai2024federated],
FedPCL [tan2022federated], FedBRB [xu2024fedbrb]
Privacy and Security Security and Robustness  [lu2023exploring],  [zhao2023privacy][wu2024vulnerabilities],  [han2023fedsecurity]
Privacy FedPIT [zhang2024fedpit], FFA-LoRA [sun2024improving]
Attacks  [wu2024vulnerabilities],  [gupta2022recovering], Decepticon [fowl2022decepticons],  [li2023unveiling]
Defense FEDML-HE [jin2023fedml],  [yu2023federated], FedBiOT [wu2023fedbiot]
Efficiency Training Efficiency Grouper [charles2024towards], FedYolo [zhang2023fedyolo], FedTune [chen2022fedtune]
Communication Efficiency CEFHRI [khalid2023cefhri], FedKSeed [qin2023federated], FedRDMA [zhang2024fedrdma]
@@
 [sun2022exploring], FedPETuning [zhang2023fedpetuning],  [wang2024personalized],
SLoRA [babakniya2023slora], FFA-LoRA [sun2023improving],
FlexLoRA [bai2024federated], LP-FL [jiang2023low], FedBiOT [wu2024fedbiot]
Frameworks Cross-Silo FedRDMA [zhang2024fedrdma], CROSSLM [deng2023mutual]
Cross-Device  [wang2023can], FwdLLM [xu2023federated],  [woisetschlager2023federated]
Decentralized Training OpenFedLLM [ye2024openfedllm],  [tang2023fusionai],  [yuan2022decentralized]
Blackbox and Transfer Learning Fed-BBPL  [lin2023efficient], ZooPFL  [lu2023zoopfl],  [kang2023grounding]
Instruction Tuning FederatedScope-LLM [kuang2023federatedscope], FedIT  [zhang2023towards]
Evaluation Datasets FederatedScope-LLM [kuang2023federatedscope], OpenFedLLM [ye2024openfedllm]
Benchmarks FedLLM-Bench  [ye2024fedllm], Fedmlsecurity [han2023fedmlsecurity], Profit [collins2023profit]
Convergence Analysis FedPEAT [chua2023fedpeat], FedMeZO [ling2024convergence]
@@ 2.1 Heterogeneity
Data Heterogeneity To address data heterogeneity for various vision-language tasks in federated multimodal fine-tuning, FedDAT [chen2023feddat] proposes a Dual-Adapter Teacher to regularize clients’ local updates and Mutual Knowledge Distillation for knowledge transfer. RaFFM [yu2023bridging] proposes a novel model compression method to accommodate heterogeneous training data. [wang2022fedkc] focuses on addressing the challenges of data heterogeneity in multilingual federated natural language understanding by proposing a FedKC module that exchanges knowledge across clients without sharing raw data. [wang2023federated] proposes a distillation method from heterogeneous tag sets to achieve better domain adaptation. In addition to transferring multimodal knowledge in FL, M2FEDSA [zhangenhancing] proposes split learning to realize modularized decomposition and enables unimodal local models to be complementarily aggregated.
Personalization and Model Heterogeneity Heterogeneity FedLoRA [yi2023fedlora] designs a homogeneous adapter to accommodate federated clients’ heterogeneous local model training, which can be aggregated on the fine-tuning server. HETLORA [cho2024heterogeneous] instead uses heterogeneous LoRA designs across devices to reduce training imbalanced problems, and sparsity-weighted aggregation is developed on the server. FlexLoRA [bai2024federated] also enables heterogeneous LoRA ranks, which are aggregated on the server using Singular Value Decomposition. FedPCL [tan2022federated] is proposed to federally fine-tune client-specific and class-relevant information with homogeneous local models as prototypes. By prototype-wise contrastive learning, the aggregated model can preserve personalized model representation while achieving similar performance with a large-scale centrally trained model. FedBRB [xu2024fedbrb] also proposes a small-to-large FL framework, where small local models are fine-tuned as blocks of a large global model.
@@ 2.2 Privacy and Security
Security and Robustness [lu2023exploring] comprehensively discussed potential vulnerabilities and essential components in decentralized training for LLMs in terms of data and models. [zhao2023privacy] studied a comprehensive literature survey about privacy-preserving in federated training of foundation models. [wu2024vulnerabilities] investigated the vulnerability of federated foundation models under adversarial threats. In addition, a novel attack strategy is developed to target the safety issues of foundation models in the setting of FL.
Privacy To enhance privacy protection and improve model performance, FedPIT [zhang2024fedpit] proposed to use LLMs’ in-context learning abilities to augment local synthetic datasets. FFA-LoRA [sun2024improving] fixes the randomly initialized non-zero matrices and only fine-tunes the zero-initialized matrices in LoRA to solve learning inefficiency in privacy-preserved FL.
Attacks Several papers also discuss the potential adversarial attack strategies and situations in federated LLMs fine-tuning. Wu et al.  [wu2024vulnerabilities] investigated the vulnerability in federated fine-tuning under adversarial threats. Experiments are conducted in both image and text domains, where this work discovered the high susceptibility of federated fine-tuning to adversarial attacks. Attackers can also recover private text in federated learning of language models [gupta2022recovering] with high fidelity. Further experiments show that if in the fine-tuning stage, the pre-trained language model’s word embedding remains unchanged, such attacks can be less effective. Decepticon [fowl2022decepticons] also proposes a latent vector attack, which can discover private user text with malicious parameter vectors. In heterogeneous federated learning, Li et al.  [li2023unveiling] also develop a novel backdoor attack mechanism by mimicking normal client behavior.
Defense Defense for privacy-preserving and malicious attacks can be beneficial and challenging for LLMs, due to the efficiency-effectiveness balance for large-scale models. FEDML-HE [jin2023fedml] proposes a homomorphic encryption algorithm that selectively encrypts sensitive parameters and balances security and efficiency. The optimized federated fine-tuning method can significantly reduce the computation overhead. [yu2023federated] more comprehensively studied the federated foundation model in privacy-preserving pre-training, fine-tuning, and application. In addition, intellectual property protection is also crucial in federated LLM fine-tuning, where FedBiOT [wu2023fedbiot] creates a bi-level optimization problem to ensure the emulator, trained on a public dataset by the LLM owner supporting adaptors in fine-tuning on clients’ private datasets, which works effectively regardless of distribution drift between the datasets.
@@ 2.3 Efficiency
Training Efficiency In LLM-based federated learning, optimizing training efficiency is crucial for minimizing latency and computational overhead. This ensures rapid and effective fine-tuning across diverse and distributed private datasets. Dataset Grouper [charles2024towards], creates large-scale federated datasets to address the data efficiency problem in federated LLM fine-tuning, which support flexible and scalable federated learning simulations for training language models with billions of parameters. FedYolo [zhang2023fedyolo] explores possibilities in on-device training of large language models, emphasizing the benefits of model scale and modularity to enhance heterogeneity robustness, reduce communication costs, and prevent catastrophic forgetting while enabling clients to adapt to diverse tasks with a single general-purpose model. FedTune [chen2022fedtune] benchmarks three fine-tuning methods (e.g., modifying the input, adding extra modules, and adjusting the backbone) and conducts empirical experiments to show the comprehensive comparison between federated learning and ordinary fine-tuning.
@@
Communication Efficiency Managing communication costs while ensuring data privacy is paramount in federated fine-tuning for large language models. Recent research has explored innovative solutions to reduce communication overhead while maintaining or enhancing the performance of federated systems. The integration of federated learning in HRI and robots benefits from data privacy but suffers from communication inefficiency. Khalid et al. [khalid2023cefhri] propose a communication-efficient FL framework for HRI (CEFHRI), leveraging pre-trained models and a trainable spatiotemporal adapter for video understanding tasks. FedKSeed [qin2023federated], explores federated full-parameter fine-tuning, which can be communication inefficient in LLMs, which utilizes zeroth-order optimization with finite random seeds, significantly reducing transmission requirements to a few thousand bytes. FedRDMA [zhang2024fedrdma], a cross-silo FL system, integrates RDMA into the FL communication protocol to enhance efficiency by dividing the updated model into chunks and optimizing RDMA-based communication.
@@
Parameter Efficiency Several existing papers have explored parameter-efficient learning methods for fine-tuning large language models (LLMs) in federated learning (FL) [wu2024fedbiot]. These methods address the communication, computation, and storage constraints of edge devices, making federated training more efficient. [sun2022exploring] systematically investigates ways to reduce communication in federated fine-tuning, while [zhang2023fedpetuning] study four PETuning methods and propose a benchmark to defend against privacy attacks. [wang2024personalized] introduce a gradient-free prompt tuning method for better performance in few-shot settings. In heterogeneous FL environments, several LoRA-based methods have been developed: SLoRA [babakniya2023slora] enables fine-tuning with sparse parameter masks and SVD-decomposed LoRA blocks, FFA-LoRA [sun2023improving] mitigates LoRA instability due to data heterogeneity, and FlexLoRA [bai2024federated] adjusts LoRA ranks dynamically. LP-FL [jiang2023low] leverages few-shot prompt learning and soft labeling for efficient communication.
@@ 2.4 Frameworks
Cross-Silo Federated learning (FL) in cross-silo settings presents unique challenges, particularly in communication efficiency and collaborative learning between different model sizes. FedRDMA [zhang2024fedrdma] is a communication-efficient FL system integrating Remote Direct Memory Access (RDMA) to alleviate the communication overheads exacerbated by LLMs. FedRDMA significantly improves communication efficiency over traditional TCP/IP-based systems by segmenting the updated model and employing optimization techniques. CROSSLM [deng2023mutual] uses smaller language models (SLMs) trained on private task-specific data to alleviate privacy concerns, which can further augment the learning of LLMs. This mutual enhancement approach allows both SLMs and LLMs to improve their performance while preserving LLMs’ generalization capabilities.
@@
Cross-Device Different from cross-silo settings, normally, cross-device federated learning involves a large number of small-scale end devices, where fine-tuning LLMs can be challenging. A systematic study [wang2023can] has used large-scale public data and LLMs to explore distillation methods in training large-scale parallel devices to improve privacy-utility tradeoff. To address the challenge of resource constraints on mobile devices, FwdLLM [xu2023federated] proposes a backpropagation-free training protocol that enhances memory and time efficiency in federated learning of LLMs, facilitating the deployment of billion-parameter LLMs on commercial mobile devices. Woisetschläger et al. [woisetschlager2023federated] focus on energy efficiency in federated learning of LLMs on edge devices. The FLAN-T5 foundation models are evaluated end-to-end for fine-tuning with a real-time metric of energy efficiency to assess the computational efficiency.
@@
Decentralized Training The extensive computational resources required for training and deploying LLMs have become a significant challenge. Recent research has explored decentralized FL as a solution, leveraging underutilized consumer-level GPUs and private data to make LLM training more accessible and privacy-preserving. OpenFedLLM [ye2024openfedllm] proposes a complete pipeline for training LLMs on the underutilized distributed private data via FL. Specifically, to leverage consumer-level GPUs for LLM training, Tang et al. [tang2023fusionai] proposed a decentralized system that includes a dynamic broker with a backup pool to handle the variability and heterogeneity of devices. Intermediate representation and execution planes are abstracted to ensure compatibility across different devices and deep learning frameworks. Yuan et al. [yuan2022decentralized] propose a scheduling algorithm that efficiently allocates computational tasks across decentralized GPUs connected by slow networks, which enables decentralized training among heterogeneous and lower-bandwidth interconnected devices.
@@
Blackbox and Transfer Learning When dealing with black-box foundation models, gradients are often unavailable, where zeroth-order optimization techniques are explored as a means to update models without requiring gradient access. Fed-BBPL  [lin2023efficient] optimizes prompt generators in federated learning settings without directly accessing the foundation models’ parameters for text and image classification tasks. ZooPFL  [lu2023zoopfl] presents a more general approach, making it suitable for more complex and diverse FL tasks. [kang2023grounding] instead focus on transfer learning that enables clients to benefit from general knowledge learned in federated settings and also transfer domain-specific knowledge back to improve the federated learning.
@@
Instruction Tuning FederatedScope-LLM [kuang2023federatedscope] includes multiple efficient federated fine-tuning algorithms to reduce communication and computation costs, where clients may or may not have access to the full model parameters. FedIT  [zhang2023towards] designs FL approaches to incorporate diverse user instructions on local devices, which might be data sensitive.